# Automating the Diagnosis of Human Vision Disorders by Cross-modal 3D Generation

### v1

@@ Abstract
Understanding the hidden mechanisms behind human’s visual perception is a fundamental quest in neuroscience, underpins a wide variety of critical applications, e.g. clinical diagnosis. To that end, investigating into the neural responses of human mind activities, such as functional Magnetic Resonance Imaging (fMRI), has been a significant research vehicle. However, analyzing fMRI signals is challenging, costly, daunting, and demanding for professional training. Despite remarkable progress in artificial intelligence (AI) based fMRI analysis, existing solutions are limited and far away from being clinically meaningful. In this context, we leap forward to demonstrate how AI can go beyond the current state of the art by decoding fMRI into visually plausible 3D visuals, enabling automatic clinical analysis of fMRI data, even without healthcare professionals. Innovationally, we reformulate the task of analyzing fMRI data as a conditional 3D scene reconstruction problem. We design a novel cross-modal 3D scene representation learning method, Brain3D, that takes as input the fMRI data of a subject who was presented with a 2D object image, and yields as output the corresponding 3D object visuals. Importantly, we show that in simulated scenarios our AI agent captures the distinct functionalities of each region of human vision system as well as their intricate interplay relationships, aligning remarkably with the established discoveries of neuroscience. Non-expert diagnosis indicate that Brain3D can successfully identify the disordered brain regions, such as V1, V2, V3, V4, and the medial temporal lobe (MTL) within the human visual system. We also present results in cross-modal 3D visual construction setting, showcasing the perception quality of our 3D scene generation.
@@ 1 Introduction
Traditional approaches in neuroscience and life sciences, such as animal studies [10, 11, 12] and invasive techniques [13, 14], have provided valuable insights but are limited by several factors. These limitations include the need for animal sacrifice [15], ethical concerns [16], and interspecies heterogeneity [17], which can hinder the translation of findings to human vision. Additionally, traditional methods often lack the spatial and temporal resolution necessary to capture the complex dynamics of neural activity underlying 3D vision perception. These limitations have restricted the progress of research in this field and highlight the need for novel interdisciplinary approaches.
Artificial vision, which aims to develop computational models inspired by biological vision [18, 19, 20], has emerged as a promising avenue for understanding the neural basis of visual perception. While artificial vision models have achieved remarkable success in various computer vision tasks, such as object recognition [21, 22, 23], detection [24, 25, 26], and scene understanding [27, 28], their contribution to understanding biological vision remains limited. These models primarily focus on replicating human-like performance on specific tasks using visual input, without directly incorporating brain signals or mimicking the neural processes involved in biological vision. Furthermore, these models are not designed to simulate the effects of brain signal perturbations on vision, which is crucial for diagnosing and understanding vision-related brain disorders.
Figure 1: Brain3D: Processing brain fMRI for clinical diagnosis. The subject’s visual system first precepts the object stimulus, triggering specific patterns of activity in the brain’s visual processing areas which is then captured and expressed by fMRI data. Brain3D then leverages a fMRI encoder to extract the feature, feeds it into a 2D diffusion model for detailed 3D generation. By contrasting the object stimulus and generated 3D visuals, even a non-expert can conduct precise diagnosis of atypical neurological conditions.
@@
Figure 2: Overview of our Brain3D that decodes the functional MRI (fMRI) signals into 3D reconstructions. Initially, the subject’s visual system precepts the object stimulus, triggering specific patterns of activity in the brain’s visual processing areas. These patterns are recorded as fMRI signals. To generate the 3D object model, Brain3D leverages two specialized encoders: the high-level encoder captures abstract high-level visual concepts, while the low-level encoder focuses on processing low-level visual details. The extracted features are passed into a two-stage reconstruction pipeline to generate the 3D representation of the object. The training procedure involves optimizing low-level and high-level encoders with the feature distillation objective $L_{image}$ and $L_{text}$ respectively, each incorporating bidirectional CLIP [72] contrastive learning and MixUp [71] augmentation for enhanced generalization.
@@ 2 Results
Overview of fMRI-based diagnosis through Brain3D.
As a bio-inspired method our Brain3D can understand how the brain interprets and processes visual stimuli in 3D, thus allowing clinical diagnosis (Fig. 2). It efficiently maps 2D visual stimuli into 3D generations by analyzing the intricate functionalities of distinct brain regions.
Specifically, our model first deploys two specialized neural encoders to encode fMRI data into multi-level neural embeddings. The low-level encoder is designed to encode fMRI data from the basic visual cortex (V1 to V4), focusing on visual details such as color, shape, and texture, into low-level neural embeddings. Concurrently, the high-level encoder concentrates on the medial temporal lobe (MTL) to process visual information, such as object semantics and types. It encodes fMRI data from the MTL regions to high-level embeddings. Additionally, we introduce a UMAP-based projection [65] to further reduce noise in the high-level embeddings by decomposing these embeddings into separate distinct representations.
Subsequently, we design a two-stage diffusion method [30, 31, 32] to reconstruct the corresponding 3D objects. 3D objects are generated through distilling a 2D diffusion model with condition on both the high level and low level fMRI embeddings, using Score Distillation Sampling (SDS) [33]. For high-fidelity visual production, we employ a coarse-to-fine optimization strategy, with the Neural Radiance Field (NeRF) [34] as 3D scene representation for initial geometry learning followed by 3D mesh representation learning for detail recovery.
With generated 3D visuals, we conduct clinical fMRI diagnosis on atypical neurological conditions. We randomly select samples from NSD testing set with simulated disorder over five brain regions: V1, V2, V3, V4, and MTL. Then, we invite 100 non-experts to detect which brain region has disorder by comparing the original image and the generated multi-view images by our model. The results affirm Brain3D’s effectiveness in identifying and diagnosing neurological abnormalities across multiple brain regions, including V1, V2, V3, V4, and the medial temporal lobe (MTL).
@@
Brain3D decode 3D models solely from fMRI signals with finer semantics in comparison with previous 2D models.
To test the effectiveness of Brain3D in 3D generation, we perform comprehensive comparisons with an existing 2D method [35]. We establish 3D-based evaluation benchmarks, including 3D Inception [36], 3D CLIP [37], 3D EfficientNet [38], and 3D Swav [39], for subsequent evaluations (see Sec. 4.6 for detailed descriptions of these benchmarks).
From the fMRI data Brain3D extracts rich appearance, semantic and geometric information, presenting a high degree of consistency with the stimuli images (Fig. 3). Despite limited observation with the giraffe, skier and teddy bear in the three cases (see the last two rows of Fig. 3), our bio-inspired model can process 3D details and comprehend their semantics beyond what is immediately visible in the 2D image, achieving finer semantic details than MindEye [35]. This result underscores the remarkable human’s neural processing capability in constructing intricate 3D structures by interpreting and internalizing 3D geometric information from 2D visual stimuli [40].
@@
Specifically, we observe that the two hemispheres make different holistic contributions to object perception, due to exhibiting distinct performances for specific objects (Fig. 4b). The left hemisphere tends to depict finer details and intricate structures in object visualization, while the right hemisphere better captures the overall shape and silhouette. These results align with previous discoveries that the left hemisphere is involved in detail-oriented tasks while the right hemisphere in more holistic tasks [43, 44]. However, our results challenge the notion of strict functional division. We found notable overlaps and interactions between the two hemispheres. Remarkably, numerous scenes rendered by the right hemisphere, as illustrated in Fig. 4a and Fig. s7, also exhibit intricate textures—typically a characteristic attributed to the left hemisphere. Conversely, the left hemisphere demonstrated capability in detecting overarching shapes, which is usually a strength of the right hemisphere.
@@
Brain3D can diagnose regions with disorder (preliminary study).
@@
The Semantic Stage within our framework is designed to extract finer details from fMRI embeddings, enhancing its contribution to 3D generation. After the Semantic Stage the detailed textures of the building in the first row and the outline of the steamship in the second row in Fig. s4c exhibit much greater clarity. These details from fMRI are challenging to represent in a coarse NeRF scene but can be effectively distilled within a 3D mesh.
@@ 3 Discussion
Understanding how human vision system extracts and processes the physical world is a critical question in neuroscience, with vital applications in healthcare and other fields. For the first time, this research models the underlying mechanisms of human visual perception by mapping fMRI of human brain activities to the corresponding 3D objects. This goes beyond previous studies that are limited to statistical fMRI analysis [42, 1, 2, 7].
Despite the advancements in our 3D scene generation and clinical diagnosis, discrepancy remains between the generated objects and the actual images viewed by participants. This phenomenon can be attributed in part to cognitive biases [78] and the inherent blurring effects [79] that occur during the image-to-fMRI cognitive stage. An example of this can be seen in the last row of Fig. 3 where the bird’s beak appears shut. This is because the upper half of the beak appears similar as the background, illustrating how visual processing and brain interpretation can influence our visual perception. Signal resolution also plays a significant role. The 1.8 mm resolution of fMRI data leads to a loss of fine details, particularly in smaller neuronal structures. In addition, current 3D generation models also struggle to accurately capture background details that extend beyond the camera’s field of view. These limitations highlight the challenges in accurately representing and analyzing complex visual information.
@@ 4.1 Pipeline overview
We aim to model the intricate mechanisms behind human vision perception. To achieve this, we develop Brain3D, a model capable of creating 3D representations from the fMRI signal of individual’s minds. Given the fMRI signal collected during an individual viewing a 2D image, we decompose it into two parts according to the sources: visual cortex and MTL. The visual cortex focuses on the low-level texture feature, while the MTL captures the high-level semantics information. For each source, we design a specialized feature embedding encoder.
Formally, this process begins with an fMRI voxel $\mathbf{X}_{i}$. We employ Multilayer Perceptrons (MLPs) as encoders to transform $\mathbf{X}_{i}$ into two types of embeddings: a low-level embedding $f_{L}(\mathbf{X}_{i})$, and a high-level embedding $f_{H}(\mathbf{X}_{i})$. These embeddings are then supervised by respective features under image and textural feature distillation. This multi-level structure thus separates the high-level semantics information and the low-level texture details of the fMRI signal, simulating the biological information processing [64, 63]. The final 3D objects are reconstructed using a two-stage generative diffusion pipeline conditioned on both the low-level and high-level fMRI embeddings.
@@ 4.2 Preliminary: Diffusion based 3D generation
This mechanism is further extended to 3D generation through the integration of Score Distillation Sampling (SDS) [33]. It involves two key components: a scene model and a pre-trained Text-to-Image (T2I) diffusion model. The scene model, parameterized by $\theta$, generates differentiable images through $\mathbf{X_{i}}=g(\theta)$. The pre-trained T2I diffusion model predicts the sampled noise $\epsilon_{\phi}(\mathbf{X_{i}}_{t};t,\mathbf{y})$ conditioned on the noisy image $\mathbf{X_{i}}_{t}$, created by adding Gaussian noise with a random scale factor $t$ to $\mathbf{X_{i}}$. The optimization of this extended model is guided by the following objective:
@@ 4.3 Tackling high noises with high-level semantics
In modeling human mind signals, a major challenge arises from the intrinsically significant levels of noise and diversity with the high-level information within the human brain (see Fig. s6). This can cause instability in the conditional probability distribution $p_{\phi}(\mathbf{z}_{t}|c,\mathbf{y})$ (see Sec. 4.2 for more details), particularly when considering different camera angles, denoted as $c$. This instability could result in a lack of consistency for 3D object generation. Addressing this issue is critical for ensuring that our model reliably interprets and reconstructs the complex fMRI scans.
We develop a novel approach to solving the above problem using the Uniform Manifold Approximation and Projection (UMAP) [65]. The key idea is to distill the high-level semantic embeddings into a series of representative embedding vectors with a unique projection weight.
Formally, it projects the high-level embedding $f_{H}(\mathbf{X}_{i})$ into a set of stable embeddings, $\left\{f_{H_{1}}(\mathbf{X}_{i}),f_{H_{2}}(\mathbf{X}_{i}),\ldots,f_{H_{n}}(\mathbf{X}_{i})\right\}$, derived from the Generative Text-to-Image Transformer [66]. In our design, we concatenate the embeddings extracted from all training images. This results in a comprehensive set of embeddings, $\left\{f_{H}(\mathbf{X}_{i}),f_{H_{1}}(\mathbf{X}_{i}),f_{H_{2}}(\mathbf{X}_{i}),\ldots,f_{H_{n}}(\mathbf{X}_{i})\right\}$, for each image. These combined embeddings form a high-dimensional dataset, denoted as D.

### v2

# Neural 3D decoding for human vision diagnosis

@@ Abstract
Understanding the hidden mechanisms behind human’s visual perception is a fundamental question in neuroscience. To that end, investigating into the neural responses of human mind activities, such as functional Magnetic Resonance Imaging (fMRI), has been a significant research vehicle. However, analyzing fMRI signals is challenging, costly, daunting, and demanding for professional training. Despite remarkable progress in artificial intelligence (AI) based fMRI analysis, existing solutions are limited and far away from being biologically meaningful and practically useful. In this context, we leap forward to demonstrate how AI can go beyond the current state of the art by advancing from 2D visuals to visually plausible and functionally more comprehensive 3D visuals decoded from brain signals, enabling automatic more sophisticated modeling of fMRI data. Innovationally, we reformulate the task of analyzing fMRI data as a conditional 3D object generation problem. We design a novel 3D object representation learning method, Brain3D, that takes as input the fMRI data of a subject who was presented with a 2D image, and yields as output the corresponding 3D object visuals. Importantly, we show that our AI agent captures the distinct functionalities of each region of human vision system as well as their intricate interplay relationships, aligning remarkably with the established discoveries of neuroscience. Non-expert diagnosis indicate that Brain3D can successfully identify the disordered brain regions in simulated scenarios, such as V1, V2, V3, V4, and the medial temporal lobe (MTL) within the human visual system. We also present results in cross-modal 3D visual generation setting, showcasing the perception quality of our 3D generation.
@@ 1 Introduction
Figure 1: Brain3D: Processing brain fMRI for clinical diagnosis. The subject’s visual system first precepts the object stimulus, triggering specific patterns of activity in the brain’s visual processing areas which is then captured and expressed by fMRI data. Brain3D then leverages a fMRI encoder to extract the feature, feeds it into a 2D diffusion model for detailed 3D generation. The training procedure involves optimizing low-level and high-level encoders with the feature distillation objective $L_{image}$ and $L_{text}$ respectively, each incorporating bidirectional CLIP [86] contrastive learning and MixUp [85] augmentation for enhanced generalization. By contrasting the object stimulus and generated 3D visuals, even a non-expert can conduct precise diagnosis of atypical neurological conditions.
Traditional approaches in neuroscience and life sciences, such as animal studies [10, 11, 12] and invasive techniques [13, 14], have provided valuable insights but are limited by several factors. These limitations include the need for animal sacrifice [15], ethical concerns [16, 17], and interspecies heterogeneity [18], which can hinder the translation of findings to human vision. Additionally, traditional methods often lack the spatial and temporal resolution necessary to capture the complex dynamics of neural activity underlying 3D vision perception. These limitations have restricted the progress of research in this field and highlight the need for novel interdisciplinary approaches.
Artificial vision, which aims to develop computational models inspired by biological vision [19, 20, 21], has emerged as a promising avenue for understanding the neural basis of visual perception. While artificial vision models have achieved remarkable success in various computer vision tasks, such as object recognition [22, 23, 24], detection [25, 26, 27], and scene understanding [28, 29], their contribution to understanding biological vision remains limited. These models primarily focus on replicating human-like performance on specific tasks using visual input, without directly incorporating brain signals or mimicking the neural processes involved in biological vision. Furthermore, these models are not designed to simulate the effects of brain signal perturbations on vision, which is crucial for understanding vision-related brain regions.
Neural decoding models have focused on extracting essential representations of brain signals for visual content prediction [30, 31, 32, 33] and object recognition [34, 35]. These investigations have also shown promise in reconstructing images [36] from fMRI data using techniques such as generative adversarial networks [37, 38] and latent-space diffusion [39, 40]. However, these studies are currently limited to single views, which constrains their ability to model and analyze the underlying human brain’s 3D visual processing capabilities.
@@ 2 Results
Overview of fMRI-based 3D sceene synthesis by Brain3D.
As a bio-inspired method our Brain3D can understand how the brain interprets and processes visual stimuli in 3D (Fig. 6). It efficiently maps 2D visual stimuli into 3D generations by analyzing the intricate functionalities of distinct brain regions.
Specifically, our model first deploys two specialized neural encoders to encode fMRI data into multi-level neural embeddings(Fig. 6a-d). The low-level encoder is designed to encode fMRI data from the basic visual cortex (V1 to V4), focusing on visual details such as color, shape, and texture, into low-level neural embeddings. Concurrently, the high-level encoder concentrates on the medial temporal lobe (MTL) to process visual information, such as object semantics and types. It encodes fMRI data from the MTL regions to high-level embeddings. Additionally, we introduce a UMAP-based projection [77] to further reduce noise in the high-level embeddings by decomposing these embeddings into separate distinct representations.
Subsequently, we design a two-stage diffusion method [42, 43, 44] to reconstruct the corresponding 3D objects. 3D objects are generated through distilling a 2D diffusion model with condition on both the high level and low level fMRI embeddings, using Score Distillation Sampling (SDS) [45](Fig. 6f). For high-fidelity visual production, we employ a coarse-to-fine optimization strategy, with the Neural Radiance Field (NeRF) [46] as 3D scene representation for initial geometry learning followed by 3D mesh representation learning for detail recovery(Fig. 6e). The results affirm Brain3D’s effectiveness in modeling human vision process and diagnosing neurological abnormalities across multiple brain regions, including V1, V2, V3, V4, and the medial temporal lobe (MTL).
@@
Brain3D decodes 3D models solely from fMRI signals with finer semantics in comparison to previous 2D models.
To test the effectiveness of Brain3D in 3D generation, we perform comprehensive comparisons with an existing 2D method [47]. We establish 3D-based evaluation benchmarks, including 3D Inception [48], 3D CLIP [49], 3D EfficientNet [50], and 3D Swav [51], for subsequent evaluations (see Sec. 4.6 for detailed descriptions of these benchmarks). From the fMRI data Brain3D extracts rich appearance, semantic and geometric information, presenting a high degree of consistency with the stimuli images (Fig. 2). Despite limited observation with the giraffe, skier and teddy bear in the three cases (see the last two rows of Fig. 2), our bio-inspired model can process 3D details and comprehend their semantics beyond what is immediately visible in the 2D image, achieving finer semantic details than MindEye [47]. This result underscores the remarkable human’s neural processing capability in constructing intricate 3D structures by interpreting and internalizing 3D geometric information from 2D visual stimuli [52].
@@
Specifically, we observe that the two hemispheres make different holistic contributions to object perception, due to exhibiting distinct performances for specific objects (Fig. 3b). The left hemisphere tends to depict finer details and intricate structures in object visualization, while the right hemisphere better captures the overall shape and silhouette. These results align with previous discoveries that the left hemisphere is involved in detail-oriented tasks while the right hemisphere in more holistic tasks [55, 56]. However, our results challenge the notion of strict functional division. We found notable overlaps and interactions between the two hemispheres. Remarkably, numerous objects rendered by the right hemisphere, as illustrated in Fig. 3a and Fig. s7, also exhibit intricate textures—typically a characteristic attributed to the left hemisphere. Conversely, the left hemisphere demonstrated capability in detecting overarching shapes, which is usually a strength of the right hemisphere.
@@
Brain3D can diagnose regions with disorder.
@@
The Semantic Stage within our framework is designed to extract finer details from fMRI embeddings, enhancing its contribution to 3D generation. After the Semantic Stage the detailed textures of the building in the first row and the outline of the steamship in the second row in Fig. s4c exhibit much greater clarity. These details from fMRI are challenging to represent in a coarse NeRF but can be effectively distilled within a 3D mesh.
@@ 3 Discussion
Understanding how human vision system extracts and processes the physical world is a critical question in neuroscience. For the first time, this research models the underlying mechanisms of human visual perception by mapping fMRI of human brain activities to the corresponding 3D objects. This goes beyond previous studies that are limited to statistical fMRI analysis [54, 1, 2, 7].
Despite the advancements in our 3D generation and clinical diagnosis, discrepancy remains between the generated objects and the actual images viewed by participants. This phenomenon can be attributed in part to cognitive biases [92] and the inherent blurring effects [93] that occur during the image-to-fMRI cognitive stage. An example of this can be seen in the last row of Fig. 2 where the bird’s beak appears shut. This is because the upper half of the beak appears similar as the background, illustrating how visual processing and brain interpretation can influence our visual perception. Signal resolution also plays a significant role. The 1.8 mm resolution of fMRI data leads to a loss of fine details, particularly in smaller neuronal structures. In addition, current 3D generation models also struggle to accurately capture background details that extend beyond the camera’s field of view. These limitations highlight the challenges in accurately representing and analyzing complex visual information.
@@ 4 Method
Figure 6: Overview of our Brain3D that decodes the functional MRI (fMRI) signals into 3D reconstructions. (a) Initially, the subject’s visual system precepts the object stimulus, which triggers specific patterns of neural activity within the brain’s visual processing areas. These patterns are recorded as fMRI signals. (b) To generate the 3D object, Brain3D employs two specialized encoders: the high-level encoder uses MLP and UMAP [77] to capture abstract, high-level visual concepts, while the low-level encoder processes basic visual details via Versatile diffusion [81]. (c) Both encoders process fMRI to extract respective information for representing 3D objects. (d) These encoders are optimized through feature distillation, which utilize pretrained CLIP-Text and CLIP-Image encoders [86] as the guidance for the high-level and low-level features, respectively. (e) The 3D object generation process includes two phases: the perceptual phase utilizes NeRF [46], and the semantic phase transforms NeRF output into a 3D mesh [83] using DMTet [82]. (f) The 3D object is refined using Score Distillation Sampling (SDS) [45], which introduces Gaussian noise into a randomly-viewed rendered image and employs 2D diffusion models for noise prediction. (g) The high-level and low-level feature guidance both are served as conditions for the pretrained Stable Diffusion [43] and Zero-1-to-3 [79] models to predict the added noise.
@@ 4.1 Pipeline overview
We aim to model the intricate mechanisms behind human vision perception. To achieve this, we develop Brain3D, a model capable of creating 3D representations from the fMRI signal of individual’s minds. Given the fMRI signal collected during an individual viewing a 2D image, we decompose it into two parts according to the sources: visual cortex and MTL(Fig. 6a). The visual cortex focuses on the low-level texture feature, while the MTL captures the high-level semantics information. For each source, we design a specialized feature embedding encoder(Fig. 6b).
Formally, this process begins with an fMRI voxel $\mathbf{X}_{i}$. We employ encoders to transform $\mathbf{X}_{i}$ into two types of embeddings: a low-level embedding $f_{L}(\mathbf{X}_{i})$, and a high-level embedding $f_{H}(\mathbf{X}_{i})$(Fig. 6c). These embeddings are then supervised by respective features under image and textural feature distillation(Fig. 6d). This multi-level structure thus separates the high-level semantics information and the low-level texture details of the fMRI signal, simulating the biological information processing [76, 75]. The final 3D objects are reconstructed using a two-stage generative diffusion pipeline conditioned on both the low-level and high-level fMRI embeddings(Fig. 6e).
@@ 4.2 Preliminary: Diffusion based 3D generation
This mechanism is further extended to 3D generation through the integration of Score Distillation Sampling (SDS) [45](Fig. 6f). It involves two key components: a scene model and a pre-trained Text-to-Image (T2I) diffusion model. The scene model, parameterized by $\theta$, generates differentiable images through $\mathbf{X_{i}}=g(\theta)$. The pre-trained T2I diffusion model predicts the sampled noise $\epsilon_{\phi}(\mathbf{X_{i}}_{t};t,\mathbf{y})$ conditioned on the noisy image $\mathbf{X_{i}}_{t}$, created by adding Gaussian noise with a random scale factor $t$ to $\mathbf{X_{i}}$. The optimization of this extended model is guided by the following objective:
@@ 4.3 Tackling high noises with high-level semantics
In modeling human mind signals, a major challenge arises from the intrinsically significant levels of noise and diversity with the high-level information within the human brain (Fig. s6). This can cause instability in the conditional probability distribution $p_{\phi}(\mathbf{z}_{t}|c,\mathbf{y})$ (see Sec. 4.2 for more details), particularly when considering different camera angles, denoted as $c$. This instability could result in a lack of consistency for 3D object generation. Addressing this issue is critical for ensuring that our model reliably interprets and reconstructs the complex fMRI scans.
We develop a novel approach to solving the above problem using the Uniform Manifold Approximation and Projection (UMAP) [77](Fig. 6b). The key idea is to distill the high-level semantic embeddings into a series of representative embedding vectors with a unique projection weight.
Formally, following the transformation of fMRI $\mathbf{X}_{i}$ into a high-level embedding by the MLP backbone $f_{H}$, this embedding $f_{H}(\mathbf{X}_{i})$ is subsequently projected
into a set of stable embeddings, $\left\{f_{H_{1}}(\mathbf{X}_{i}),f_{H_{2}}(\mathbf{X}_{i}),\ldots,f_{H_{n}}(\mathbf{X}_{i})\right\}$, derived from the Generative Text-to-Image Transformer [78].
In our design, we concatenate the embeddings extracted from all training images. This results in a comprehensive set of embeddings, $\left\{f_{H}(\mathbf{X}_{i}),f_{H_{1}}(\mathbf{X}_{i}),f_{H_{2}}(\mathbf{X}_{i}),\ldots,f_{H_{n}}(\mathbf{X}_{i})\right\}$, for each image. These combined embeddings form a high-dimensional dataset, denoted as D.